# Are you a robot? Detecting Autonomous Vehicles from Behavior Analysis

### v1

@@ I Introduction
With the advent of emerging communication technologies, the automotive sector has exhibited relevant industrial profits towards the so-called revolution of connected and autonomous cars. This digital transformation enables advanced services drawing attention to two crucial societal aspects, such as safety of road users and sustainability of involved infrastructure, thereby delaying the transition to connected roads with fully-autonomous cars as part of the SAE level 3+: this implies vehicles implementing an environmental perception module [1]. Within this transition phase, it appears clear a need for the coexistence of human drivers and autonomously-driven vehicles: computers will take over the maneuvering control of the cars to take specific actions (e.g., highway driving or parking), while giving it back upon critical situations [2].
@@
Such mixed traffic conditions [3] will involve a number of severe safety issues that must be tackled [4]: autonomously-driven vehicles will slowly build the risk perception [5] exacerbating safeguarding concerns; it is crucial to spot autonomous cars with (potential) inconsistent behaviors [6]. We propose an overarching framework to automatically detect whether a vehicle is autonomous or human-driven based on gathered behavioral features and to provide feedbacks for improving autonomous systems: it benchmarks established machine learning models trained on temporal series of images and additional numeric features as the state information of the target vehicle to classify.
To tackle this problem, we built and exploited a dataset that collects human and autonomous driving scenes, in the form of videos, i.e., sequences of RGB images. These are recorded by the EGO vehicle (EV) following a target vehicle (TV), whose state, collected alongside, is characterized by speed, acceleration, rotation, distance from the EV and distance from lane center. We have built the dataset using the CARLA simulator [7] and the Baidu Apollo autonomous driving agent to control the target in the autonomous settings [8, 9], whereas the human driving behavior has been acquired by manually steering the TV with a Logitech G29 steering wheel. The dataset is dubbed as NexusStreets [^1] NexusStreets is inspired by the Nexus series of replicants in Blade Runner, who were designed to be nearly indistinguishable from humans. The name emphasizes the idea of a mix of human-driven and autonomous vehicles on the same streets. and is publicly released on Zenodo at https://zenodo.org/records/7682484.
@@ I-A Related Work
A useful taxonomic characterization of behavior prediction methods is based on the type of input features employed by the models [10]. In [11, 12], the authors propose machine learning models that predict vehicle behavior from the track history of the TVs states. The accuracy of such methods is limited by the performance of the EV’s perception module. Note that they do not consider vehicle-vehicle, nor vehicle-environment interaction patterns. In [13] the authors propose a deep learning framework to automatically learn regular mobile traffic patterns along roads thereby detecting non-recurring events and classify them by severity level to assist the emergency operations and support advanced services provided by the network operator.
When other surrounding vehicles’ track histories are considered, patterns are extracted from the interactions between them and the target vehicles [14, 15]. Analogously to the works presented above, these methods’ performance is limited by the accuracy of the EV’s perception that provides the track histories. However, they can learn from a richer input representations and hence capture more complex dynamics.
@@
To account for vehicle-environment interactions, a further class of methods inputs simplified bird’s eyes views of the scene [16, 17]. Finally, [18, 19] propose to directly predict vehicle behavior from the raw sensor data to overcome the information loss injected by the perception stack.
@@ I-B Contributions
The paper contributions can be summarized as follows: $i$) we generate a novel dataset combining autonomous and human driving scenes, $ii$) we present benchmark results of various machine learning models on the driving scene classification task, $iii$) we model the problem as a supervised classification of multivariate time series describing the temporal evolution of the Target Vehicle (TV) state, including a second multivariate time series representing 2D bounding box detection of the TV in the pixel coordinates space over time, $iv$) we carry out an exhaustive simulation campaign to validate our framework, $v$) we make the NexusStreets dataset publicly available to the scientific community. Fig. 2 provides a schematic illustration of the full pipeline.
Our experiments show that it is possible to discriminate with high accuracy ($>0.95$ auROC) whether a TV is autonomous or human-driven alone from the observation of its driving behavior in the form of state and 2D object detection.
@@ II-A Simulation Environment and Considered Scenarios
Autonomous and manual driving scenes have been collected using a nearly-realistic simulator, namely CARLA [7] used in driving-related research. It offers access to the vehicles’ sensors and formally describes the roads according to OpenDRIVE specifications [20]. It also provides interfaces with external tools and software, e.g., ROS [21, 22], and autonomous driving agents to control vehicles. Additionally, it allows the integration of external steering wheels to directly drive the vehicles: human behaviors have been reproduced with a Logitech G29 steering wheel, which supports realistic force feedback and first-person driving experience.
@@ II-B Description of the Main Actors
The dataset properly combines autonomous driving scenes wherein the TV is implemented using the Baidu Apollo [8, 9] autonomous driving agent, and human-driven scenes wherein the TV is controlled by skilled people directly driving with a steering wheel in a first-person perspective [23].
@@ II-C Sampling Scenes and Target States
In addition to the TV state, we provide 2D object detections in the form of bounding box coordinates w.r.t the reference system $F$ at each timestamp (see Fig. 4). We operate 2D object detection using a pre-trained YOLOv7 [24] model. This convolutional neural network (CNN) outputs bounding boxes and class probabilities for each object detected in the image. For each acquired frame, we consider objects labeled as “cars” with a confidence $\geq 0.9$ and ignore other classes. We store the detection as vectors $\bm{x^{D}}=\big{[}x,y,h,w\big{]}$, where $(x,y)$ are the bounding box center coordinates, $h$ its height and $w$ its width. For each frame, multiple vehicles are detected. It is therefore necessary to identify the bounding box of the TV. To achieve this goal, for each weather and light condition, we first collect a set of $2$ images (as template) depicting the TV rear and side. Then, we embed the template images using a ResNet-50v2 [25] model trained on ImageNet [26]. Embedding vectors are obtained from the activation of the final pre-softmax fully-connected layer. Afterwards, for each frame, we embed all detection crops and then compute their average Euclidean distance from the TV template. The detection crop whose embedding presents the smallest average Euclidean distance is designated as the true TV detection [^3] Analogously, to a real-world perception module, this approach is not exempted from noise. Nevertheless, the combination of YOLO and ResNet-50v2-based template matching is able to detect the TV most of the times. Mistakes are considered outliers given their small numbers..
@@ III-A Vehicle Behavior Classification
Machine Learning Models. We benchmark two different machine learning models, namely a Random Forest (RF) [27] and a deep neural network based on Long Short-Term Memory (LSTM) cells [28].
Figure 7: Vehicle behavior classification results. These have been obtained by applying RF and LSTM models to the datasets $\mathcal{X}^{S}$ and $\mathcal{X}^{D}$ separately. Reported values are test mean scores over $5$ repeated experiments. Confidence intervals are standard deviations.
@@
[…]
@@
Sliding Windows Sampling. The samples contained in the datasets $\mathcal{X}^{S}$, $\mathcal{X}^{D}$ and $\mathcal{X}^{S+D}$ are constituted by the full history of states and 2D detection associated to the scenes, which span $60$s and present $120$ timestamps each. Given that our aim is to make predictions based on shorter time intervals, we sample time windows of shorter duration from each scene. First, we define $5$ values for time window duration: $5$s, $10$s, $15$s, $20$s and $25$s. Then, for each window size, we apply a sliding window approach and extract sub-scenes from the whole dataset of the defined time duration. We train the machine learning models on the obtained sub-scenes. This allows to deeply observe how the scene duration can affect the models performance.
@@
Figure 8: Effect of state information loss across various drop rates. The degradation has been simulated by obfuscating state information from the $X^{S+D}$ dataset with various drop rates. $r=0$ corresponds to the best case scenario where all the data is intact. Test scores are averaged over $5$ repeated experiments.
@@
Figure 9: Effect of different history and future horizon lengths. Test results for future state autoregression are averaged over $5$ repeated experiments. Error bars represent standard deviation. RMSE: root mean squared error. Time duration is quantified in terms of timestamp number. $1$ timestamp is approximately $0.5$s.
@@ III-B Future State Autoregression
In the following, we present experimental results for the future state autoregression task. The term “auto” in autoregressive indicates that a variable is regressed against itself [29]. We first present a formalization of the problem. Then, we show the effect of different past (i.e., history) and future horizon lengths on the prediction error. Finally, we compare the performance achieved by the model on scenes wherein the target vehicle (TV) is controlled by the autonomous agent against scenes in which the vehicle is driven by a human.
@@
Figure 10: Future state autoregression error for autonomous agent and human driver. Error bars represent the standard deviation over $5$ repeated experiments with different training/test splits.
@@ IV Risk-based Training Process
An additional result from the comparison described in the previous section is that the lower predictability of the autonomous behavior can result in an increased exposure to the safety risks, as it leads to abrupt braking and/or sudden acceleration or deceleration. This is usually considered among the major causes of hazards for the bystanders and the EV passengers [30], as an exhaustive hazards analysis and risk assessment conducted according to ISO 26262 [31] could highlight. The higher predictability of human driving could result from the inherent human capability to perceive the risk in diverse scenarios [32], which suggests the idea to perform an iterative training process of the autonomous driving control algorithm using as objective function the difference between autonomous and human driver prediction errors or similar metrics based on risk estimation, to finally minimize unpredictable behavior effects with subsequent safety risks. This iterative training process is out of the scope of this paper and will be the subject of future works. Future works will also seek to integrate data sourced from real vehicles, validating its suitability for real world applications.
@@ V Conclusions
In this paper, we have presented a novel autonomous vehicle detection solution capable of identifying whether vehicles are driven by humans or computers. We generated and released publicly a custom dataset for this purpose, NexusStreet. It combines both autonomous-driven scenes, obtained on the CARLA simulator by means of Baidu Apollo control agent, and human-driven ones, acquired by using a steering wheel maneuvered by licensed (and skilled) human drivers. After training various machine learning models, we show that it is possible to automatically detect autonomous vehicles by evaluating video segments. Accuracy and auROC improves to $+90\%$ when combined with histories of the state vector, i.e., position, speed, acceleration and orientation, locally estimated by all installed sensors. Moreover, information as state vector and local classification can be shared among road users through Vehicle-to-everything (V2X)-based infrastructure, further increasing the overall autonomous vehicles detection accuracy to fine-tune the car autonomous control system.

### v2

@@ I Introduction
With the advent of emerging communication technologies, the automotive sector has exhibited relevant industrial profits towards the so-called revolution of connected and autonomous cars. This digital transformation enables advanced services drawing attention to two crucial societal aspects, such as safety of road users and sustainability of involved infrastructure, thereby delaying the transition to connected roads with fully-autonomous cars as part of the SAE level 3+: this implies vehicles implementing an environmental perception module [duarte2018impact]. Within this transition phase, it appears clear a need for the coexistence of human drivers and autonomously-driven vehicles: computers will take over the maneuvering control of the cars to take specific actions (e.g., highway driving or parking), while giving it back upon critical situations [virdi2019safety].
@@
Such mixed traffic conditions [parth2017] will involve a number of severe safety issues that must be tackled [9636795]: autonomously-driven vehicles will slowly build the risk perception [peng2023sotif] exacerbating safeguarding concerns; it is crucial to spot autonomous cars with (potential) inconsistent behaviors [hussain2022deepguard]. We propose an overarching framework to automatically detect whether a vehicle is autonomous or human-driven based on gathered behavioral features and to provide feedbacks for improving autonomous systems: it benchmarks established machine learning models trained on temporal series of images and additional numeric features as the state information of the target vehicle to classify.
To tackle this problem, we built and exploited a dataset that collects human and autonomous driving scenes, in the form of videos, i.e., sequences of RGB images. These are recorded by the EGO vehicle (EV) following a target vehicle (TV), whose state, collected alongside, is characterized by speed, acceleration, rotation, distance from the EV and distance from lane center. We have built the dataset using the CARLA simulator [Dosovitskiy17] and the Baidu Apollo autonomous driving agent to control the target in the autonomous settings [huang2018apolloscape, ap], whereas the human driving behavior has been acquired by manually steering the TV with a Logitech G29 steering wheel. The dataset is dubbed as NexusStreets [^1] NexusStreets is inspired by the Nexus series of replicants in Blade Runner, who were designed to be nearly indistinguishable from humans. The name emphasizes the idea of a mix of human-driven and autonomous vehicles on the same streets. and is publicly released on Zenodo at https://zenodo.org/records/7682484.
@@ I-A Related Work
A useful taxonomic characterization of behavior prediction methods is based on the type of input features employed by the models [mozaffari2020deep]. In [zyner2019naturalistic, xin2018intention], the authors propose machine learning models that predict vehicle behavior from the track history of the TVs states. The accuracy of such methods is limited by the performance of the EV’s perception module. Note that they do not consider vehicle-vehicle, nor vehicle-environment interaction patterns. In [piroad_infocom21] the authors propose a deep learning framework to automatically learn regular mobile traffic patterns along roads thereby detecting non-recurring events and classify them by severity level to assist the emergency operations and support advanced services provided by the network operator.
When other surrounding vehicles’ track histories are considered, patterns are extracted from the interactions between them and the target vehicles [diehl2019graph, ma2019trafficpredict]. Analogously to the works presented above, these methods’ performance is limited by the accuracy of the EV’s perception that provides the track histories. However, they can learn from a richer input representations and hence capture more complex dynamics.
@@
To account for vehicle-environment interactions, a further class of methods inputs simplified bird’s eyes views of the scene [lee2017convolution, schreiber2019long]. Finally, [luo2018fast, casas2018intentnet] propose to directly predict vehicle behavior from the raw sensor data to overcome the information loss injected by the perception stack.
@@ I-B Contributions
The paper contributions can be summarized as follows: $i$) we generate a novel dataset combining autonomous and human driving scenes, $ii$) we present benchmark results of various machine learning models on the driving scene classification task, iiiiii) we model the problem as a supervised classification of multivariate time series describing the temporal evolution of the Target Vehicle (TV) state, including a second multivariate time series representing 2D bounding box detection of the TV in the pixel coordinates space over time, $iv$) we carry out an exhaustive simulation campaign to validate our framework, $v$) we make the NexusStreets dataset publicly available to the scientific community. Fig. 2 provides a schematic illustration of the full pipeline.
Our experiments show that it is possible to discriminate with high accuracy (>0.95>0.95 auROC) whether a TV is autonomous or human-driven alone from the observation of its driving behavior in the form of state and 2D object detection.
@@ II-A Simulation Environment and Considered Scenarios
Autonomous and manual driving scenes have been collected using a nearly-realistic simulator, namely CARLA [Dosovitskiy17] used in driving-related research. It offers access to the vehicles’ sensors and formally describes the roads according to OpenDRIVE specifications [dupuis2010opendrive]. It also provides interfaces with external tools and software, e.g., ROS [quigley2009ros, macenski2022robot], and autonomous driving agents to control vehicles. Additionally, it allows the integration of external steering wheels to directly drive the vehicles: human behaviors have been reproduced with a Logitech G29 steering wheel, which supports realistic force feedback and first-person driving experience.
@@ II-B Description of the Main Actors
The dataset properly combines autonomous driving scenes wherein the TV is implemented using the Baidu Apollo [huang2018apolloscape, ap] autonomous driving agent, and human-driven scenes wherein the TV is controlled by skilled people directly driving with a steering wheel in a first-person perspective [haris2022navigating].
@@ II-C Sampling Scenes and Target States
In addition to the TV state, we provide 2D object detections in the form of bounding box coordinates w.r.t the reference system $F$ at each timestamp (see Fig. 4). We operate 2D object detection using a pre-trained YOLOv7 [wang2023yolov7] model. This convolutional neural network (CNN) outputs bounding boxes and class probabilities for each object detected in the image. For each acquired frame, we consider objects labeled as “cars” with a confidence $\geq 0.9$ and ignore other classes. We store the detection as vectors $\bm{x^{D}}=\big[x,y,h,w\big]$, where $(x,y)$ are the bounding box center coordinates, $h$ its height and $w$ its width. For each frame, multiple vehicles are detected. It is therefore necessary to identify the bounding box of the TV. To achieve this goal, for each weather and light condition, we first collect a set of $2$ images (as template) depicting the TV rear and side. Then, we embed the template images using a ResNet-50v2 [he2016deep] model trained on ImageNet [deng2009imagenet]. Embedding vectors are obtained from the activation of the final pre-softmax fully-connected layer. Afterwards, for each frame, we embed all detection crops and then compute their average Euclidean distance from the TV template. The detection crop whose embedding presents the smallest average Euclidean distance is designated as the true TV detection [^3] Analogously, to a real-world perception module, this approach is not exempted from noise. Nevertheless, the combination of YOLO and ResNet-50v2-based template matching is able to detect the TV most of the times. Mistakes are considered outliers given their small numbers..
@@ III-A Vehicle Behavior Classification
Machine Learning Models. We benchmark two different machine learning models, namely a Random Forest (RF) [ho1995random] and a deep neural network based on Long Short-Term Memory (LSTM) cells [hochreiter1997long].
Figure 7: Vehicle behavior classification results. These have been obtained by applying RF and LSTM models to the datasets $\mathcal{X}^{S}$ and $\mathcal{X}^{D}$ separately. Reported values are test mean scores over 55 repeated experiments. Confidence intervals are standard deviations.
@@
For the RF, we optimize the number of trees over $\{100k\>|\>1\leq k\leq 10\}$, the minimum number of samples required to be at a leaf node over $\{k\>|\>1\leq k\leq 5\}$, the function to measure the quality of a split over $\{\text{Gini impurity},\text{Shannon entropy}\}$ and the number of features to consider when looking for the best split over $\{\sqrt{n},\log_{2}n\}$, where $n$ is the number of input features. Since the RF inputs vectors, we unroll the input matrices $\bm{X}$ to 11D vectors.
@@
Sliding Windows Sampling. The samples contained in the datasets $\mathcal{X}^{S}$, $\mathcal{X}^{D}$ and $\mathcal{X}^{S+D}$ are constituted by the full history of states and 2D detection associated to the scenes, which span $60$s and present $120$ timestamps each. Given that our aim is to make predictions based on shorter time intervals, we sample time windows of shorter duration from each scene. First, we define $5$ values for time window duration: $5$s, $10$s, $15$s, 2020s and $25$s. Then, for each window size, we apply a sliding window approach and extract sub-scenes from the whole dataset of the defined time duration. We train the machine learning models on the obtained sub-scenes. This allows to deeply observe how the scene duration can affect the models performance.
@@
Figure 8: Effect of state information loss across various drop rates. The degradation has been simulated by obfuscating state information from the $X^{S+D}$ dataset with various drop rates. $r=0$ corresponds to the best case scenario where all the data is intact. Test scores are averaged over 55 repeated experiments.
@@
Figure 9: Effect of different history and future horizon lengths. Test results for future state autoregression are averaged over 55 repeated experiments. Error bars represent standard deviation. RMSE: root mean squared error. Time duration is quantified in terms of timestamp number. 11 timestamp is approximately $0.5$s.
@@ III-B Future State Autoregression
In the following, we present experimental results for the future state autoregression task. The term “auto” in autoregressive indicates that a variable is regressed against itself [hyndman2018forecasting]. We first present a formalization of the problem. Then, we show the effect of different past (i.e., history) and future horizon lengths on the prediction error. Finally, we compare the performance achieved by the model on scenes wherein the target vehicle (TV) is controlled by the autonomous agent against scenes in which the vehicle is driven by a human.
@@
Figure 10: Future state autoregression error for autonomous agent and human driver. Error bars represent the standard deviation over 55 repeated experiments with different training/test splits.
@@ IV Risk-based Training Process
An additional result from the comparison described in the previous section is that the lower predictability of the autonomous behavior can result in an increased exposure to the safety risks, as it leads to abrupt braking and/or sudden acceleration or deceleration. This is usually considered among the major causes of hazards for the bystanders and the EV passengers [IVS2017], as an exhaustive hazards analysis and risk assessment conducted according to ISO 26262 [iso26262] could highlight. The higher predictability of human driving could result from the inherent human capability to perceive the risk in diverse scenarios [Kolekar2020], which suggests the idea to perform an iterative training process of the autonomous driving control algorithm using as objective function the difference between autonomous and human driver prediction errors or similar metrics based on risk estimation, to finally minimize unpredictable behavior effects with subsequent safety risks. This iterative training process is out of the scope of this paper and will be the subject of future works. Future works will also seek to integrate data sourced from real vehicles, validating its suitability for real world applications.
@@ V Conclusions
In this paper, we have presented a novel autonomous vehicle detection solution capable of identifying whether vehicles are driven by humans or computers. We generated and released publicly a custom dataset for this purpose, NexusStreet. It combines both autonomous-driven scenes, obtained on the CARLA simulator by means of Baidu Apollo control agent, and human-driven ones, acquired by using a steering wheel maneuvered by licensed (and skilled) human drivers. After training various machine learning models, we show that it is possible to automatically detect autonomous vehicles by evaluating video segments. Accuracy and auROC improves to $90\%$ when combined with histories of the state vector, i.e., position, speed, acceleration and orientation, locally estimated by all installed sensors. Moreover, information as state vector and local classification can be shared among road users through Vehicle-to-everything (V2X)-based infrastructure, further increasing the overall autonomous vehicles detection accuracy to fine-tune the car autonomous control system.